OpenStudy (anonymous):

y=mx+b form Slope-3; through (-3,-2) a) y=-3x+11 b) y=-3x-11

OpenStudy (anonymous):

b

OpenStudy (aroub):

y=mx+b , m=-3 , (-3,-2) -2=-3(-3)+b -2=9+b -11=b =>y=-3x-11 (b)